OpenStudy (anonymous):

Can anyone find the tangent line approximation for f(x)=x^4+3x^2, when a =-1?

myininaya (myininaya):

f '(x)=4x^3+6x did you get this far?

OpenStudy (anonymous):

ok i know how you got there

myininaya (myininaya):

f '(1)=4+6=10

myininaya (myininaya):

10 is the slope of the tangent line at x=-1

myininaya (myininaya):

so we know a point on the line and we know the slope of that line and a line has form y=mx+b f(-1)=10(-1)+b we can find b now 1+3=-10+b 4=-10+b 10+4=b 14=b so the tangent line is y=10x+14

OpenStudy (anonymous):

For whatever reason the answers in the back of the book are telling me the tangent would equal -10x-6, is it possible we mixed signs along the way?

myininaya (myininaya):

let me do it on paper one sec

myininaya (myininaya):

is this what you are confused about? the answer in the back of the book? did you try this problem and get what I got?

myininaya (myininaya):

oh i plug in 1

OpenStudy (anonymous):

You didn't plug in -1

myininaya (myininaya):

right polpak

OpenStudy (anonymous):

No i got something like x^5+5x^4+7x^3+9x^2+6x because i thought it was F(x)=f(a)+f'(a)(x-a) inorder to solve this beast

myininaya (myininaya):

f '(-1)=-10 not 10

myininaya (myininaya):

y=-10x+b we need to know b we know a point on this line

OpenStudy (anonymous):

Nah, just find the slope of the tangent and plug in that slope into point slope formula

OpenStudy (anonymous):

Ok after you get the slope of the tangent what did you plug in to get you (x,y) values?

myininaya (myininaya):

If x=1, then y=....

myininaya (myininaya):

damn stupid 1

myininaya (myininaya):

-1

OpenStudy (anonymous):

\(x = -1 \implies f = x^4+3x^2 = ?\)

myininaya (myininaya):

lol

OpenStudy (anonymous):

thats 4 but we are stillnot jiving with the stupid books answer

myininaya (myininaya):

we are fixing to

myininaya (myininaya):

so we have y=-10x+b 4=-10(-1)+b

OpenStudy (anonymous):

Take the slope (-10) and the point (-1,4) and plug into point slope. \(y-y_0 = m(x-x_0)\)

myininaya (myininaya):

solve for b

myininaya (myininaya):

you can use either way

OpenStudy (anonymous):

Where m is the slope, \(x_0,y_0\) are the x and y for your point.

myininaya (myininaya):

b=4-10=-6 so y=-10x-6

OpenStudy (anonymous):

I like point slope cause I don't have to do anything else. Slope intercept is only useful if you're gonna graph it.

myininaya (myininaya):

i like slope intercept its more fun to me for some reason

OpenStudy (anonymous):

Ok where did we pull the y value from (4) sorry i missed how we came up with that one

OpenStudy (anonymous):

Plug in x=-1 to the original equation for f(x)

OpenStudy (anonymous):

Nevermind I'm sorry, I just figured that part out, thanks for the help guys. It has been a long day to put it nicely.